OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@Loser66

OpenStudy (nurali):

g(x) = 2(x − 4) 20= 2(x − 4) 20/2=x-4 10=x-4 x=10+4 x=_______

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

Correct